OpenStudy (anonymous):

*fan and medal awarded* What is the value of the y variable in the solution to the following system of equations? 5x + 3y = 7 3x - 5y = -23 -1 1 -4 4

OpenStudy (anonymous):

-4 is my answer

OpenStudy (anonymous):

\[5x + 3y = 7\]\[ 3x - 5y = -23\] multiply the first equation all the way across by 3, and the second by \(-5\)

OpenStudy (anonymous):

oops no you are off by a minus sign

OpenStudy (anonymous):

oh my bad

OpenStudy (anonymous):

Auditions for a play in New York were held with 100 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 20, how many men auditioned? 44 40 76 64 i searched around for this one and got 76 is that right ?

OpenStudy (anonymous):

but im thinkin 64 now.... idk help !

OpenStudy (anonymous):

\[15x+9y=21\] \[-15x+25y=115\] \[34y=136\] \[y=136\div 34=4\]

OpenStudy (anonymous):

thanks ! can u help with the other ? i got 64 off of basic common sence

OpenStudy (anonymous):

yes that is right

OpenStudy (anonymous):

cool i didnt even do real math lol

OpenStudy (anonymous):

common sense is good

OpenStudy (anonymous):

sure you did, it is called "thinking"

OpenStudy (anonymous):

yeaa but im thinking it over now and it doesnt make any sence

OpenStudy (anonymous):

ok we can use an equation if you like

OpenStudy (anonymous):

but i didnt use equations or anything. theres a difference between math and common sence math. yes please

OpenStudy (anonymous):

if you put \(x\) as the number of men, and \(y\) as the number of women, then since the total is 100 you know \(x+y=100\)

OpenStudy (anonymous):

and the second weird sentence tells you that \(2x-3y=20\)

OpenStudy (anonymous):

so you can solve \[x+y=100\]\[2x-3y=20\] in a couple different ways one way is to rewrite the first one as \(y=100-x\) and then substitute in to the second equation to get \[2x-3(100-x)=20\]

OpenStudy (anonymous):

multiply out to get \[2x-300+3x=20\] so \[5x-300=20\] making \[5x=320\] so \[x=320\div 5=64\]

OpenStudy (anonymous):

and the 300 came from the 3(100 - x) right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well actually it is \(-300\)

OpenStudy (anonymous):

oh okay thanks !

OpenStudy (anonymous):

yw